# Uniform convergence of the smooth calibration error and its relationship with functional gradient

### v2

@@ C.4 Proof of Corollary 1
[…]
@@
$|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|\leq 24\sqrt{2}C\,\mathfrak{R}_{\mathcal{D},2n}(\mathcal{F})(\log 2n)^{2}+\frac{32}{\sqrt{2n}}+\sqrt{\frac{\log(1/\delta)}{n}}.$ (153)
C.5 Proof of Theorem 2
@@
They further showed that there exists a universal constant $C$ such that
based on the result of ambroladze2004complexity.
We now allocate the total failure probability $\delta$ across the two terms using the union bound. Specifically, we set $\delta\to\frac{2}{3}\delta$ in Eq. (154), and $\delta\to\frac{1}{3}\delta$ in Theorem 1 or Corollary 1. Then, the sum of the confidence terms becomes:
$\sqrt{\frac{\log\frac{2}{\frac{2}{3}\delta}}{2n}}+\sqrt{\frac{\log\frac{1}{\frac{1}{3}\delta}}{n}}=\left(1+\frac{1}{\sqrt{2}}\right)\sqrt{\frac{\log\frac{3}{\delta}}{n}}\leq\sqrt{\frac{2\log\frac{3}{\delta}}{n}}.$ (157)
Thus, we obtain the desired high-probability bound. ∎
C.6 Additional remarks about the smooth CE
@@
Under this setting, for any $f\in\mathcal{F}$, there exists $\tilde{f}$ such that
$|\mathrm{smCE}(f,\mathcal{D},S_{n})-\mathrm{smCE}(\tilde{f},\mathcal{D},S_{n})|\leq\epsilon$ (158)
Proof.
Then define the corresponding SMCE as
$\mathrm{smCE}(f,\mathcal{D},S_{n})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\omega^{*}_{i}$ (159)
and we also have
$\mathrm{smCE}(\tilde{f},\mathcal{D},S_{n})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\tilde{v}_{i})\tilde{\omega}^{*}_{i}$ (160)
. Then we have
$\mathrm{smCE}(f,\mathcal{D},S_{n})-\mathrm{smCE}(\tilde{f},\mathcal{D},S_{n})$ $=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\omega^{*}_{i}-\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\tilde{\omega}^{*}_{i}$ (161)
$\leq\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\omega^{*}_{i}-\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i}){\omega}^{*}_{i}$ (162)
$=\frac{1}{n}\sum_{i=1}^{n}(\tilde{v}_{i}-v_{i})\omega^{*}_{i}$ (163)
$\leq\sqrt{\frac{1}{n}\sum_{i=1}^{n}|f(X_{i})-f^{\prime}(X_{i})|^{2}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}|\omega^{*}_{i}|^{2}}$ (164)
$\leq\sqrt{\frac{1}{n}\sum_{i=1}^{n}|f(X_{i})-f^{\prime}(X_{i})|^{2}}\leq\epsilon$ (165)
where the first inequality is due to the definition of $\tilde{\omega}^{*}$ and the second inequality is due to the Cauchy-Schwartz inequality.
As for the other direction, we have
$\mathrm{smCE}(f,\mathcal{D},S_{n})-\mathrm{smCE}(\tilde{f},\mathcal{D},S_{n})$ $=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\omega^{*}_{i}-\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\tilde{\omega}^{*}_{i}$ (166)
$\geq\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\tilde{\omega}^{*}_{i}-\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\tilde{\omega}^{*}_{i}$ (167)
$=\frac{1}{n}\sum_{i=1}^{n}(\tilde{v}_{i}-v_{i})\tilde{\omega}^{*}_{i}$ (168)
$\geq-\sqrt{\frac{1}{n}\sum_{i=1}^{n}|f(X_{i})-f^{\prime}(X_{i})|^{2}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}|\tilde{\omega}^{*}_{i}|^{2}}$ (169)
$\geq-\sqrt{\frac{1}{n}\sum_{i=1}^{n}|f(X_{i})-f^{\prime}(X_{i})|^{2}}\leq-\epsilon$ (170)
In conclusion, we have
$|\mathrm{smCE}(f,\mathcal{D},S_{n})-\mathrm{smCE}(\tilde{f},\mathcal{D},S_{n})|\leq\epsilon$ (171)
Appendix D Proofs of the gradient boosting tree
D.1 Proof of Theorem 3
@@
Recall that the step size is set to $w_{t}=1/(t+1)$. Let $\nabla_{t,i}=\nabla_{g}\ell(\sigma(g^{(t)}(X_{i})),Y_{i})$ denote the functional gradient at the $i$-th training point.
Algorithm 1 L1-Constrained Gradient Boosting Tree with Cross-Entropy Loss
@@
5:     Solve: $\theta_{t}=\operatorname*{\mathop{\rm argmin}}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\psi_{\theta}(X_{i})\nabla_{t,i}$
@@
8:  return  $g^{(T)}(x)$
Proof.
By the definition of the supremum, there exists $\eta^{*}$ such that
$\sup_{\eta}\frac{1}{n}\sum_{i=1}^{n}\left[-\nabla_{t,i}\eta(g^{(t)}(X_{i}))\right]\leq\frac{1}{n}\sum_{i=1}^{n}\left[-\nabla_{t,i}\eta^{*}(g^{(t)}(X_{i}))\right]+\epsilon.$ (172)
where ϵ+italic-$\epsilon\in\mathbb{R}^{+}$ is any positive constant.
Consider a function $f(x)$ expressed as a convex combination of base learners:
$f(x)=\sum_{m=1}^{2M}u_{m}\psi_{\tilde{\theta}_{m}}(x),$ (173)
where $u_{m}\geq 0$ and $\sum_{m=1}^{2M}u_{m}=1$. Let $\Psi$ denote the set of all such convex combinations. We will specify an appropriate choice of $f$ later.
By definition of the minimizer at each step, for any $\psi_{\tilde{\theta}_{m}}\in\mathcal{T}$, the following holds:
$w_{t}\sum_{i=1}^{n}\nabla_{t,i}\psi_{\theta_{t}}(X_{i})\leq w_{t}\sum_{i=1}^{n}\nabla_{t,i}\psi_{\tilde{\theta}_{m}}(X_{i}).$ (174)
Multiplying both sides by $u_{m}$ and summing over $m=1,\dots,2M$, we obtain:
$w_{t}\sum_{i=1}^{n}\nabla_{t,i}\psi_{\theta_{t}}(X_{i})\leq w_{t}\sum_{i=1}^{n}\nabla_{t,i}f(X_{i}),$ (175)
where we used the definition of $f\in\Psi$.
Next, we show that we can choose $f=\eta^{*}+g^{(t)}$. This is justified as follows: first, note that $g^{(t)}$ is a convex combination of trees, expressed as
$g^{(t)}=\sum_{t=0}^{T-1}w_{t}\prod_{t^{\prime}=t+1}^{T-1}(1-w_{t^{\prime}})\psi_{\theta_{t}},$
and since $\sum_{t=0}^{T-1}w_{t}\prod_{t^{\prime}=t+1}^{T-1}(1-w_{t^{\prime}})\leq 1/2$ by the definition of the step size, we have $g^{(t)}\in\Psi$. Therefore, there exists a set of trees such that
$\nabla_{t,i}g^{(t)}(X_{i})=\nabla_{t,i}\sum_{m=1}^{M}u_{m}\psi_{\tilde{\theta}_{m}}(X_{i})+\epsilon_{1},$ (176)
with $\sum_{m=1}^{M}u_{m}\leq 1/2$ and arbitrarily small $\epsilon_{1}$.
For $\eta^{*}$, since we assume $m\geq d$, it is possible to approximate indicator functions using suitable combinations of binary trees. Specifically, for each input $X_{i}$, we can construct an indicator function matching $\eta^{*}(g^{(t)}(X_{i}))$; see breiman2004population for details. Since we assume $B\geq 2$, we can approximate
$\nabla_{t,i}\eta^{*}(g^{(t)}(X_{i}))=\nabla_{t,i}\sum_{m=M+1}^{2M}u_{m}\psi_{\tilde{\theta}_{m}}(X_{i})+\epsilon_{2},$ (177)
with $\sum_{m=1}^{2M}u_{m}=1$ and arbitrarily small $\epsilon_{2}$.
Combining the above, for each $i\in\{1,\dots,n\}$, there exists $f_{i}\in\Psi$ such that
$\nabla_{t,i}(g^{(t)}(X_{i})+\eta^{*}(g^{(t)}(X_{i})))=\nabla_{t,i}f_{i}(X_{i})+\epsilon_{1}+\epsilon_{2}.$ (178)
Summing over $i$ and averaging, we obtain:
$\frac{1}{n}\sum_{i=1}^{n}\nabla_{t,i}(g^{(t)}(X_{i})+\eta^{*}(g^{(t)}(X_{i})))=\frac{1}{n}\sum_{i=1}^{n}\nabla_{t,i}f_{i}(X_{i})+\epsilon_{1}+\epsilon_{2}.$ (179)
Since each $f_{i}\in\Psi$, combining with Eq. (175), we get:
$\frac{1}{n}\sum_{i=1}^{n}\nabla_{t,i}\psi_{\theta_{t}}(X_{i})\leq\frac{1}{n}\sum_{i=1}^{n}\nabla_{t,i}(g^{(t)}(X_{i})+\eta^{*}(g^{(t)}(X_{i})))-\epsilon_{1}-\epsilon_{2}.$ (180)
Rearranging terms gives:
$-\frac{1}{n}\sum_{i=1}^{n}\nabla_{t,i}\eta^{*}(g^{(t)}(X_{i}))\leq\frac{1}{n}\sum_{i=1}^{n}\nabla_{t,i}(g^{(t)}(X_{i})-\psi_{\theta_{t}}(X_{i}))-\epsilon_{1}-\epsilon_{2}.$ (181)
We now apply the following lemma:
Lemma 5.
Assume $T\geq 2$, $w_{t}=1/(t+2)$, and $\sup_{\theta}\|\psi_{\theta}(\cdot)\|_{\infty}\leq B$ for some $B\geq 2$. Then,
$\min_{t\in\{0,\dots,T\}}\frac{1}{n}\sum_{i=1}^{n}\nabla_{t,i}(g^{(t)}(X_{i})-\psi_{\theta_{t}}(X_{i}))\leq\frac{6B^{2}}{T+2}.$ (182)
The proof of this lemma is given in Appendix D.2.
Letting $\epsilon,\epsilon_{1},\epsilon_{2}\to 0$, we conclude:
$\min_{t\in\{0,\dots,T\}}\sup_{\eta}\frac{1}{n}\sum_{i=1}^{n}\left[-\nabla_{t,i}\eta(g^{(t)}(X_{i}))\right]\leq\frac{6B^{2}}{T+2}.$ (183)
D.2 Proofs of Lemma 5
This lemma is adapted from Theorem 2 in jaggi2013revisiting. Although the step sizes used in our setting differ slightly from theirs, the proof of their Theorem 2 (see Appendix B) applies directly to the $\ell_{1}$-constrained Gradient Boosted Trees (GBT) in our case. Below, we highlight the key difference in step sizes under their notation.
In their Appendix B, when using their notation, modifying the step size leads to the update:
$h^{(k+1)}<h^{(k)}-\frac{2\beta C-C/\mu^{2}}{2D^{2}},$ (184)
where a new factor of 2 appears in the denominator. As a result, we need to set $\mu=1/3$ and $\beta=6$ in their proof. This modification leads to a change in the constant, and their analysis yields:
$g^{(k)}\leq\frac{\beta C}{D},$ (185)
in their notation, where $C=2C_{f}$. From their Appendix D, they show that:
$C_{f}\leq\mathrm{dism}_{\|\cdot\|}(\mathcal{D})^{2}L,$
where $L$ is the Lipschitz constant of the derivative of the loss function.
In our case, $C_{f}$ is upper bounded by the squared diameter of the domain and the Lipschitz constant $L$. Since each regression tree is uniformly bounded by $B$, the diameter is $2B$, and we have $L=1/4$. Therefore, we obtain:
$C_{f}\leq B^{2}.$
Moreover, the parameter $\delta$ appearing in their theorem is set to $\delta=0$ in our setting. Recall that $D=k+2$, thus, by applying Theorem 2 from jaggi2013revisiting with these constants, we obtain the desired result.
D.3 Proof of Corollary 2
@@
We now derive an upper bound on the Rademacher complexity. Using its sub-additive property, we have:
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})\leq\left(\sum_{t=0}^{T-1}w_{t}\right)\mathfrak{R}_{\mathcal{D},n}(\mathcal{T})\leq\mathfrak{R}_{\mathcal{D},n}(\mathcal{T}),$ (187)
@@
[…]
@@
$\ln N(\epsilon,\mathcal{T},L_{2}(S_{n}))\leq\ln(nd)^{2^{J}}\left(\frac{C}{\epsilon^{2}}\right)^{2^{J+1}}.$ (189)
@@
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{T})\leq C\sqrt{\frac{2^{J}\log(nd)}{n}}.$ (190)
Finally, substituting Eq. (190) into Theorem 2, we obtain the result.
D.4 Misclassification rate
We use the proof of the conditional gradient. Theorem 1 in jaggi2013revisiting or Theorem 3.8 in bubeck2015convex can be adapted to our setting. Similar to those proofs, we show the result by induction, then we have for $t=0,\dots,T$
$L_{n}(g^{(t)})-L_{n}(g^{*})\leq\frac{L(2B)^{2}}{t+2}=\frac{B^{2}}{t+2}$ (191)
where $g^{*}=\operatorname*{\mathop{\rm argmin}}_{g\in\Psi}L_{n}(g)$ where $\Psi$ is the class of the convex combination of trees. From the standard theory of the margin classification accuracy, for any $\gamma>0$ with probability at least $1-\delta$, we have
$P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq P_{(X,Y)\sim\mathcal{D}_{n}}[(2Y-1)\bar{g}^{(T)}(X)\leq\rho]+\frac{2}{\rho}\mathfrak{R}_{\mathcal{D},n}(\mathcal{G})+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}$ (192)
and from the property of the cross entropy loss and the logsitc loss, we have
$P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq\frac{1}{\log(1+e^{-\rho})}\Big{(}L_{n}(g^{*})+\frac{B^{2}}{t+2}\Big{)}+\frac{2}{\rho}\mathfrak{R}_{\mathcal{D},n}(\mathcal{G})+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}$ (193)
and substituting the estimate of the Rademacher complexity of Eq. (190), we get the misclassification rate bound. Note that this bound does not require the depth assumption, which is required for the smooth CE analysis.
Appendix E Proofs for the kernel boosting
E.1 Proof of Eq. (22)

### v4

@@ D.2 Sub-optimal result of the Rademacher complexity
$|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ $\leq\inf_{\epsilon>0}\left[8\epsilon+24\int_{\epsilon}^{1}\left(\frac{2\sqrt{2}\mathfrak{G}_{\mathcal{D},2n}(\mathcal{F})}{\epsilon^{\prime}}+\frac{1}{\sqrt{n}}\right)d\epsilon^{\prime}\right]+\sqrt{\frac{\log(1/\delta)}{n}}$ (191)
@@
$|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|\leq 24\sqrt{2}C\mathfrak{R}_{\mathcal{D},2n}(\mathcal{F})(\log 2n)^{2}+\frac{32}{\sqrt{2n}}+2\sqrt{\frac{\log(1/\delta)}{n}}.$ (195)
Appendix E Proofs of the gradient boosting tree
E.1 Quadratic upper bound of the boosting
The iterative minimization problem is given as
$\min_{\psi_{\theta}}\mathcal{L}_{n}\big(g^{(t)}-w_{t}\psi_{\theta}(x)\big),$ (196)
and by considering the quadratic upper bound for this
$\min_{\psi_{\theta}}\Big(-\langle\nabla\mathcal{L}_{n}(g^{(t)}),w_{t}\psi_{\theta}\rangle+\frac{M}{2}\|w_{t}\psi_{\theta}\|^{2}\Big)\Leftrightarrow\min_{\psi_{\theta}}\|Mw_{t}\psi_{\theta}-\nabla\|^{2}.$ (197)
E.2 From weak learnability to functional gradient
We use the empirical inner product and norms
$\langle a,b\rangle_{L_{2}(S_{n})}:=\frac{1}{n}\sum_{i=1}^{n}a_{i}b_{i},\quad\|f\|_{L_{2}(S_{n})}^{2}:=\langle f,f\rangle_{L_{2}(S_{n})},\quad\|v\|_{L_{1}(S_{n})}:=\frac{1}{n}\sum_{i=1}^{n}|v_{i}|.$
Consider the binary cross-entropy with logit $z=g(X)$, i.e.,
$\ell_{\rm ent}(\tilde{y},z)\ =\ -\tilde{y}\log\sigma(z)\ -\ (1-\tilde{y})\log\bigl(1-\sigma(z)\bigr),\quad\sigma(z)=\frac{1}{1+e^{-z}}.$
Its per-sample gradient w.r.t. the logit is
$\nabla_{i}\ :=\ \frac{\partial}{\partial z}\ell_{\rm ent}(\tilde{y}_{i},z)\Big|_{z=g(X_{i})}\ =\ \sigma\bigl(g(X_{i})\bigr)\ -\ \tilde{y}_{i}.$
Let $v=(\nabla_{1},\ldots,\nabla_{n})$.
Lemma 4.
Under the same assumptions in Theorem 3, the following holds:
$\mathrm{smCE}^{\sigma}\big(g^{(t)},S_{n}\big)\leq\|v\|_{L_{1}(S_{n})}\leq\frac{1}{\gamma B}\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}+\frac{Mw_{t}}{2\gamma}B.$ (198)
@@ E.2 From weak learnability to functional gradient
Introduce $y_{i}:=2\tilde{y}_{i}-1\in\{\pm 1\}$. Then
$\nabla_{i}=\sigma\bigl(g(X_{i})\bigr)-\tilde{y}_{i}=-y_{i}\sigma\bigl(-y_{i}g(X_{i})\bigr):=-y_{i}q_{i},\quad q_{i}\in(0,1),$
hence $|\nabla_{i}|=q_{i}$. Let $w_{i}:=\dfrac{|\nabla_{i}|}{\sum_{j=1}^{n}|\nabla_{j}|}\in\Delta_{n}$. By Assumption 2 there exists $\psi^{\star}\in\Psi$ such that $\sum_{i=1}^{n}w_{i}y_{i}\psi^{\star}(X_{i})\geq\gamma B$. Therefore
$\big\langle v,\psi^{\star}\big\rangle_{L_{2}(S_{n})}=\frac{1}{n}\sum_{i=1}^{n}(-y_{i}|\nabla_{i}|)\psi^{\star}(X_{i})=-\frac{1}{n}\Big(\sum_{i=1}^{n}|\nabla_{i}|\Big)\Big(\sum_{i=1}^{n}w_{i}y_{i}\psi^{\star}(X_{i})\Big)\ \leq\ -\gamma B\|v\|_{L_{1}(S_{n})}.$
Since $\Psi$ is sign-flip closed, $\psi^{\star}\in\Psi$, and
$\big\langle v,-\psi^{\star}\big\rangle_{L_{2}(S_{n})}\geq\ \gamma B\|v\|_{L_{1}(S_{n})}.$ (199)
Note that
$\|Mw_{t}\psi_{\theta_{t}}-\nabla\|_{L_{2}(S_{n})}^{2}\leq\|Mw_{t}(-\psi^{\star})-\nabla_{t}\|_{L_{2}(S_{n})}^{2},$ (200)
which implies that
$-2Mw_{t}\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}\leq-2Mw_{t}\big\langle(-\psi^{\star}),\ \nabla_{t}\big\rangle_{L_{2}(S_{n})}+M^{2}w_{t}^{2}\|\psi^{\star}\|^{2}_{L_{2}(S_{n})},$ (201)
and we obtain
$\langle(-\psi^{\star}),\ \nabla_{t}\big\rangle_{L_{2}(S_{n})}\leq\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}+\frac{Mw_{t}}{2}B^{2}.$ (202)
Then combining Eq. equation 199 and equation 202 we obtain
$\gamma B\|v\|_{L_{1}(S_{n})}\leq\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}+\frac{Mw_{t}}{2}B^{2}.$ (203)
By definition, we have
$\mathrm{smCE}^{\sigma}\big(g^{(t)},S_{n}\big)\leq\|v\|_{L_{1}(S_{n})}.$ (204)
Since $M=1/4$, combining the above two inequalities, we obtain the result. ∎
E.3 Proof of Theorem 3
We consider the Taylor expansion of the objective
$L_{n}(g^{(t+1)})$ $\leq L_{n}(g^{(t)})-w_{t}\langle\nabla,\psi_{\theta_{t}}\rangle_{L_{2}(S_{n})}+\frac{M}{2}w_{t}^{2}\|\psi_{\theta_{t}}\|_{L_{2}(S_{n})}^{2}.$ (205)
Here we use Lemma 5 in Appendix E.4. The lemma is the Pythagorean relation of the projection of the function gradient. In the lemma, $w_{t}M\to w$ and $\nabla\to x$, and $p\to\psi_{\theta_{t}}$ corresponding to the GBT step. From Eq. equation 213, setting $q=0$ and , we get
$\langle\psi_{\theta_{t}},w_{t}M\psi_{\theta_{t}}-\nabla_{t}\rangle_{L_{2}(S_{n})}\leq 0,$ (206)
and this leads to
$Mw_{t}\|\psi_{\theta_{t}}\|_{n}^{2}\leq\langle\psi_{\theta_{t}},\nabla_{t}\rangle_{L_{2}(S_{n})}.$ (207)
Finally we have
$L_{n}(g^{(t+1)})$ $\leq L_{n}(g^{(t)})-\frac{w_{t}}{2}\langle\nabla_{t},\psi_{\theta_{t}}\rangle_{L_{2}(S_{n})}.$ (208)
By summing up from 0 to T1T-1, we have
$\frac{1}{2}\sum_{t=0}^{T-1}w_{t}\langle\nabla_{t},\psi_{\theta_{t}}\rangle_{L_{2}(S_{n})}$ $\leq L_{n}(g^{(0)})-L_{n}(g^{(T)}).$ (209)
If we use the constant stepsize $w_{t}=w$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\langle\nabla_{t},\psi_{\theta_{t}}\rangle_{L_{2}(S_{n})}$ $\leq\frac{2\big(L_{n}(g^{(0)})-L_{n}(g^{(T)})\big)}{wT}.$ (210)
Combining Lemma 4, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathrm{smCE}^{\sigma}\big(g^{(t)},S_{n}\big)\leq\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla_{t}\|_{L_{1}(S_{n})}$ $\leq\frac{1}{2\gamma B}\cdot\frac{1}{T}\sum_{t=0}^{T-1}\langle\nabla_{t},\psi_{\theta_{t}}\rangle_{L_{2}(S_{n})}+\frac{wMB}{2\gamma}$ (211)
$\leq\frac{1}{2\gamma B}\cdot\frac{2\big(L_{n}(g^{(0)})-L_{n}(g^{(T)})\big)}{wT}+\frac{wMB}{2\gamma}.$ (212)
Since the cross entropy loss is always positive, we drop $L_{n}(g^{(T)})$ and obtain the result.
E.4 Auxiliary lemmas
Setting.
Let $X\subset\mathbb{R}^{d}$ be a nonempty closed convex set and fix a scalar weight $w>0$. Consider
$f(z)\ :=\ \tfrac{1}{2}\|wz-x\|_{2}^{2}\ =\ \tfrac{w^{2}}{2}\big\|z-\tfrac{x}{w}\big\|_{2}^{2}\ +\ \mathrm{const},\qquad x\in\mathbb{R}^{d},$
and let
[…]
Equivalently, with $y_{0}:=x/w$, $p$ is the Euclidean projection $p=\Pi_{X}(y_{0})$.
Following is the alternative lemma of Lemma 3.1 in bubeck2015convex,
@@ E.4 Auxiliary lemmas
For all $q\in X$, the following relation holds.
$\langle p-q,w^{2}p-wx\rangle\leq 0.$
Equivalently, since $w^{2}p-wx=w^{2}(p-y_{0})$, we have
$\langle p-q,p-y_{0}\rangle\leq 0.$ (213)
Proof. The objective $f$ is convex and differentiable with gradient
$\nabla f(z)\ =\ w^{2}z-wx.$
For convex differentiable optimization over a convex set $X$, the necessary and sufficient first-order condition at a minimizer $p$ is
$\langle\nabla f(p),q-p\rangle\geq 0\qquad(\forall q\in X).$
Rewriting gives $\langle\nabla f(p),p-q\rangle\leq 0$, i.e. $\langle p-q,w^{2}p-wx\rangle\leq 0$. $\square$
E.5 Summary of the gradient boosting tree algorithm
@@
and $|C_{j}|$ is bounded by $B$
Let $\nabla_{t,i}=\nabla_{g}\ell(\sigma(g^{(t)}(X_{i})),Y_{i})$ denote the functional gradient at the $i$-th training point.
Algorithm 1 Gradient Boosting Tree with Cross-Entropy Loss
@@
5:  Solve: $\theta_{t}=\operatorname*{arg\,min}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}|Mw_{t}\psi_{\theta}(X_{i})-\nabla_{t,i}|^{2}$
@@
8:return $g^{(T)}(x)$ or $\bar{g}^{(T)}(x)$
E.6 Proof of Corollary 2
@@
We now evaluate the Rademacher complexity. We express the function class of $\bar{g}^{(T)}=\frac{1}{n}\sum_{t=0}^{T-1}g^{t}$ as $\mathcal{G}$.
Since $\mathcal{F}=\sigma\circ\mathcal{G}$ and $\sigma$ is $1/4$-Lipschitz, Talagrand’s contraction lemma yields:
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})=\mathfrak{R}_{\mathcal{D},n}(\sigma\circ\mathcal{G})\leq\frac{1}{4}\mathfrak{R}_{\mathcal{D},n}(\mathcal{G}).$ (215)
We now derive an upper bound on the Rademacher complexity of $\mathcal{G}$. Using its sub-additive property, we have:
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{G})\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathfrak{R}_{\mathcal{D},n}(\{g{(t)}\}).$ (216)
where $\{g{(t)}\}$ is the function class that after $t$-step GBT algorithms.
Then
$\mathfrak{R}_{\mathcal{D},n}(\{g^{(t)}\})\leq wt\mathfrak{R}_{\mathcal{D},n}(\mathcal{T}),$ (217)
@@
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{T})\leq\inf_{\epsilon>0}\left[4\epsilon+12\int_{\epsilon}^{1}\sqrt{\frac{\ln N(\epsilon^{\prime},\mathcal{T},L_{2}(S_{n}))}{n}}d\epsilon^{\prime}\right].$ (218)
@@
$\ln N(\epsilon,\mathcal{T},L_{2}(S_{n}))\leq\ln(nd)^{2^{m}}\left(\frac{C}{\epsilon^{2}}\right)^{2^{m+1}}.$ (219)
@@
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{T})\leq C\sqrt{\frac{2^{m}\log(nd)}{n}}.$ (220)
In conclusion, we have that
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})\leq CwT\sqrt{\frac{2^{m}\log(nd)}{n}}.$ (221)
Finally, substituting this into Theorem 2, we obtain the result.
$\mathrm{smCE}^{\sigma}\big(\frac{1}{T}\sum_{t=0}^{T-1}g^{(t)},S_{n}\big)\leq$ $\frac{L_{n}(g^{(0)})}{\gamma BwT}+\frac{wMB}{2\gamma}+wTC\sqrt{\frac{2^{m}\log(nd)}{n}}\!+\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (222)
E.7 Misclassification rate
Combining the above results, we obtain:
$P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq(1+e^{\rho})\left(\frac{L_{n}(g^{(0)})}{\gamma BwT}+\frac{wB}{8\gamma}\right)+\frac{CwT}{\rho}\sqrt{\frac{2^{J}\log(nd)}{n}}+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}.$ (225)
This result illustrates a trade-off with respect to $wT$, which balances the training misclassification rate and the complexity of the hypothesis class.
Appendix F Proofs for the kernel boosting
F.1 Proof of Eq. (23)